# Sparse Autoencoders Trained on the Same Data Learn Different Features

### v1

@@ 2 Related work
Recent work has found that SAE features are not atomic, in the sense that a “meta SAE” can decompose them into more specific features (Anonymous, 2024). Relatedly, a feature in a small SAE may be replaced by multiple, more specific features in a larger SAE. In some cases, a more general feature like starts with the letter L appears alongside a specific feature like the token “lion”, which may prevent the general feature from being active in contexts where intuitively, both the general and the specific feature apply (Chanin et al., 2024). In light of these phenomena, some have questioned whether the “flat” design of standard SAEs can accommodate the hierarchical structure inherent to human concepts (Ayonrinde et al., 2024). While sparse autoencoders presuppose that neural networks use linear representations, some research suggests that irreducibly nonlinear features also exist (Engels et al., 2024). If this is true, SAEs trained with different random initializations might converge to different ways of “linearizing” the nonlinear features of activation space.
@@ 3 Methods
In order to measure the degree of alignment between independently trained SAEs, we use the Hungarian algorithm to efficiently compute the matching between their latents which maximizes the average cosine similarity between the matched encoder and decoder vectors. This average cosine similarity after matching can be used as an overall alignment score.
@@ 4 Asymptotic trend
The number of latents found in only one SAE decreases slowly as the number of seeds increases (Figure 2, right panel). If we could train infinite seeds the number of latents “shared” by only 1 SAE would go to zero, but our results seems to indicate that when training a small number of SAEs a certain number of “orphan” latents are always present– we found that a power law with an offset term fits the data significantly better than one without the offset.
@@ 7 Conclusion
Code to perform the Hungarian alignment of SAE seeds as well as all the analysis done in this work can be found here, and all the trained SAE checkpoints can be found here.
Acknowledgements
Gonçalo Paulo had the idea of investigating the overlap in features learned by SAEs trained with different random seeds, trained the SAEs, and ran all the experiments. Gonçalo wrote the first draft, and Nora Belrose extensively edited the draft into a final version.
Gonçalo and Nora are funded by a grant from Open Philanthropy. We thank Coreweave for computing resources.

### v2

@@ 2 Related Work
Recent work has found that SAE features are not atomic, in the sense that a “meta SAE” can decompose them into more specific features (Anonymous, 2024). Relatedly, a feature in a small SAE may be replaced by multiple, more specific features in a larger SAE. In some cases, a more general feature like starts with the letter L appears alongside a specific feature like the token “lion”, which may prevent the general feature from being active in contexts where intuitively, both the general and the specific feature apply (Chanin et al., 2024). In light of these phenomena, some have questioned whether the “flat” design of standard SAEs can accommodate the hierarchical structure inherent to human concepts (Ayonrinde et al., 2024). While sparse autoencoders presuppose that neural networks use linear representations, some research suggests that irreducibly nonlinear features also exist (Engels et al., 2024). If this is true, SAEs trained with different random initializations might converge to different ways of “linearizing” the nonlinear features in activation space.
@@
Concurrently with our work, Balagansky et al. (2024) use the Hungarian algorithm to align features from SAEs trained on adjacent layers of the Gemma 2 models (Lieberum et al., 2024). However, we learned in personal communication with the DeepMind interpretability team that the same random seed was used to initialize every SAE in the Gemmascope collection, so Balagansky et al. (2024)’s positive results are likely dependent on this hyperparameter choice.
@@ 3 Methods
In order to measure the degree of alignment between independently trained SAEs, we use the Hungarian algorithm to efficiently compute the matching between their latents which maximizes the average cosine similarity between the matched encoder and decoder vectors. This average cosine similarity after matching is the overall alignment score.
@@ 4 Asymptotic Trend
The number of latents found in only one SAE decreases slowly as the number of seeds increases (Figure 2, right panel). Our results seems to indicate that when training a small number of SAEs a certain number of “orphan” latents are always present – we found that a power law with an offset term fits the data significantly better than one without the offset.
@@ 7 Conclusion
Code to perform the Hungarian aligment of SAE seeds can be found here.